OpenStudy (anonymous):

Find all the values of x in the interval [0, 2pie] that satisfy the equation 2cos(x) + sin(2x) = 0.

OpenStudy (anonymous):

step one is to rewrite \[sin(2x)=2sin(x)cos(x)\]

OpenStudy (anonymous):

then you get \[2cos(x)+2cos(x)sin(x)=0\] \[2cos(x)(1+sin(x))=0\] \[cos(x)=0\] or \[1+sin(x)=0\] \[sin(x)=-\frac{1}{2}\]

OpenStudy (anonymous):

can you solve from there?

OpenStudy (anonymous):

can you continue please?

OpenStudy (anonymous):

sure. you are in the interval \[(0,2\pi)\]

OpenStudy (anonymous):

in that interval cosine is 0 at \[\frac{\pi}{2}\] and \[\frac{3\pi}{2}\]

OpenStudy (anonymous):

are those fractions?

OpenStudy (anonymous):

if you do not instantly know where \[sin(x)=-\frac{1}{2}\] then look at the cheat sheet http://tutorial.math.lamar.edu/cheat_table.aspx and see that it is at \[\frac{7\pi}{6}\] and \[\frac{11\pi}{6}\]

OpenStudy (anonymous):

that is where the second coordinate is \[-\frac{1}{2}\]

OpenStudy (anonymous):

unit circle on last page of cheat sheet